# Simplifying Hypergraph Neural Networks

### v2

@@ 6.1 Experiment Setup
Dataset and baseline. For the hypergraph node classification task, we conduct experiments on six real-world hypergraphs: Cora-CA, DBLP-CA, Citeseer, Congress, House, and Senate. We use nine HNNs as baselines: HGNN (Feng et al., 2019), HCHA (Bai et al., 2021), HNHN (Dong et al., 2020), UniGCNII (Huang & Yang, 2021), AllDeepSets (Chien et al., 2022), AllSetTransformer (Chien et al., 2022), PhenomNN (Wang et al., 2023b), ED-HNN (Wang et al., 2023a), and Deep-HGNN (Chen et al., 2022). For the hyperlink prediction task, we compare our SHNN with AllSetTransformer, PhenomNN, ED-HNN, and Deep-HGNN on three real-world hypergraphs: Cora-CA, Citeseer, and House. The dataset statistics are summarized in Table 3, with detailed descriptions of datasets and baselines available in Appendix.
Implementation. For hypergraph node classification, following previous works (Wang et al., 2023a, b; Duta et al., 2023), we use a 50%/25%/25% train/validation/test data split and adapt the baseline classification accuracy from these studies [^2] Since there are no reported results for Deep-HGNN under the chosen data split, we use our own results.. Additionally, similar to these works, we implement the classifier based on MLP for our SHNN and report the classification accuracy from ten runs. For hyperlink prediction, existing works (Chen & Liu, 2023) primarily focus on the design of the prediction head, with no reported results for applying our baseline HNNs to this task; so we report all results conducted by ourselves. We employ the deep set function implemented by (Chien et al., 2022) as the prediction head for our method and baseline methods due to its simplicity. We use a 50%/25%/25% train/validation/test data split and ensure each split contains five times as many fake hyperedges as real hyperedges. Moreover, both SHNN and HNNs only use real hyperedges in training and validation sets to do message passing. The learning frameworks are evaluated by Area Under the ROC curve (AUC). All experiments were performed on an RTX 3090 using PyTorch, and our code with running commands will be at https://anonymous.4open.science/r/SHNN. More details about the model hyperparameters are provided in Appendix.
@@ 6.2 Comparison with Baselines
Table 4: Node classification accuracy (%) for HNN-based and SHNN-based learning frameworks on benchmarks. The state-of-the-art result on each dataset is highlighted in bold font. The second and third highest accuracies are marked with an underline for easy identification.
@@
Figure 2: Comparison of the training time required for HNN-based and SHNN-based learning frameworks to achieve optimal performance on benchmarks.
@@
Figure 3: Hyperlink prediction Area Under the ROC curve (AUC) (%) and relative training time for HNN-based and SHNN-based learning frameworks on benchmarks.
Node classification. We summarize the classification accuracy and relative training time of the SHNN-based and HNN-based frameworks in Table 1 and Figure 2, respectively. Table 4 demonstrates that SHNN not only leads to the best results across four datasets (Cora-CA, Congress, House, and Senate) but also results in the highest average mean accuracy overall. These findings confirm the effectiveness of SHNN in generating powerful node features for node classification. Further, as shown in Figure 2, the SHNN-based framework consistently needs less training time to achieve superior performance compared to the HNN-based framework. This underscores the efficacy of SHNN in enhancing training efficiency for learning frameworks in hypergraph-related downstream tasks.
Hyperlink prediction. The results for the hyperlink prediction task are summarized in Figure 3. These results exhibit a pattern similar to that observed in the hypergraph node classification task: the SHNN-based framework outperforms the HNN-based frameworks while requiring less training time. An interesting observation is that the training efficiency improvement brought by SHNN is smaller in the hyperlink prediction task compared to the node classification task. We attribute this to the fact that the task-specific model used for hyperlink prediction is more computationally expensive than the one used for hypergraph node classification. Specifically, the hyperlink prediction model needs to generate a prediction for each hyperedge by aggregating the features of the nodes, which has a complexity of $\mathcal{O}(nd_{v})$, whereas the node classification model only needs to perform forward propagation for each individual node, with a complexity of $\mathcal{O}(n)$. Here $n$ is the number of nodes and $d_{v}$ is the average node degree. Appendix theoretically demonstrates that the efficiency improvement provided by SHNN is inversely correlated with the complexity of the task-specific model.
@@ 6.3 Ablation study
We conduct a series of ablation studies on the SHNN-based learning framework in node classification on the Cora-CA, Citeseer, House, and Senate datasets.
@@
Figure 4: The ablation study on the number of layers and the $\alpha$ of SHNN in node classification.
Table 5: The ablation study on the SHNN-based node classification framework. The state-of-the-art result on each dataset is highlighted in bold font.
SHNN Weighted $\mathbf{S}$ Cora-CA Citeseer Senate House Avg. Mean
@@
Ablation study on hyperparameters. We focus on two key hyperparameters of SHNN: the number of layers and the $\alpha$. As illustrated in Figure 4, models with $\alpha>0$ exhibit more consistent accuracy as the number of layers increases compared to models with $\alpha=0$. We attribute this consistency to the robustness discussed in Proposition 4.2, where a positive $\alpha$ helps retain the distinct information of each node, thereby mitigating the oversmoothing issue. Additionally, we observe that a positive $\alpha$ significantly enhances performance in the House and Senate datasets, which are more heterophilic, compared to the Cora-CA and Citeseer datasets. This suggests that retaining the distinct information of each node is more beneficial for heterophilic datasets.
Ablation study on learning framework components. In this study, we examine the influence of SHNN and hyperedge-size-based clique expansion on the performance of an SHNN-based node classification framework. Without SHNN, the framework functions as a pure MLP. We define $\mathbf{S}$ based on hyperedge-size-based clique expansion as weighted $\mathbf{S}$; in contrast, for the framework without weighted $\mathbf{S}$, $\mathbf{S}$ is generated using a $\mathbf{W}$ with all positive values set to one. As shown in Table 5, the inclusion of SHNN improves classification accuracy across all datasets, demonstrating that SHNN effectively incorporates hypergraph structural information into node features. Furthermore, adding weighted $\mathbf{S}$ further enhances performance, highlighting the informativeness of the hyperedge-size-based clique expansion. Thus, the components we designed significantly aid the task-specific model in leveraging useful information from hypergraph data.
@@ 7 Conclusion & Limitation
In this paper, we propose the Simplified Hypergraph Neural Network (SHNN), a novel training-free model designed specifically for hypergraph-structured data. SHNN addresses the challenge of balancing efficiency and effectiveness in the design of hypergraph machine learning models. We present both theoretical and empirical evidence demonstrating that SHNN can solve hypergraph-related downstream tasks with remarkable efficiency and effectiveness. Our SHNN can advance the field of hypergraph machine learning research by largely improving the efficiency of using observed hypergraph structures to solve downstream tasks.
Limitation and future work. Learning tasks on hypergraphs can generally be defined at the node, hyperedge, and hypergraph levels. In this paper, we empirically test our SHNN on node-level and hyperedge-level tasks, with the main limitation being the lack of empirical results on hypergraph-level tasks. To our knowledge, designing task-specific models to generate hypergraph-level features based on existing node features remains a challenging topic in the literature. Therefore, we consider the application of our SHNN to hypergraph-level tasks as future work.

### v3

@@ 6.1 Experiment Setup
Dataset and baseline. For the hypergraph node classification task, we conduct experiments on six real-world hypergraphs: Cora-CA, DBLP-CA, Citeseer, Congress, House, and Senate. We use nine HNNs as baselines: HGNN (Feng et al., 2019), HCHA (Bai et al., 2021), HNHN (Dong et al., 2020), UniGCNII (Huang & Yang, 2021), AllDeepSets (Chien et al., 2022), AllSetTransformer (Chien et al., 2022), PhenomNN (Wang et al., 2023b), ED-HNN (Wang et al., 2023a), and Deep-HGNN (Chen et al., 2022). For the hyperlink prediction task, we compare our SHNN with AllSetTransformer, PhenomNN, ED-HNN, and Deep-HGNN on three real-world hypergraphs: Cora-CA, Citeseer, and House. The dataset statistics are summarized in Table 2, with detailed descriptions of datasets available in Appendix.
Metric. Following previous works (Wang et al., 2023a; Chen & Liu, 2023), we evaluate models for node classification and hyperlink prediction with classification accuracy and the Area Under the ROC Curve (AUC), respectively. We evaluate the efficiency of the models by the relative training time $r_{f}$, defined as $r_{f}=t_{f}/t_{S}$, where $t_{f}$ is the training time to achieve optimal performance for the evaluated model, and $t_{s}$ is the training time for optimal performance of our SHNN.
Implementation. For node classification, we follow previous works (Wang et al., 2023a; Duta et al., 2023) to use a 50%/25%/25% train/validation/test data split and adapt the baseline classification accuracy from them [^3] Since there are no reported results for Deep-HGNN under the chosen data split, we use our own results.. Additionally, similar to these works, we implement the classifier based on MLP for our SHNN and report the results from ten runs. For hyperlink prediction, existing works (Chen & Liu, 2023) primarily focus on the design of the prediction head, with no reported results for applying our baseline HNNs to this task; so we report all results conducted by ourselves. We employ the deep set function implemented by (Chien et al., 2022) as the prediction head for our method and baseline methods due to its simplicity. We use a 50%/25%/25% train/validation/test data split and ensure each split contains five times as many fake hyperedges as real hyperedges. Both SHNN and HNNs only use real hyperedges in training and validation sets to do message passing. All experiments were performed on an RTX 3090 with PyTorch, and our code with running commands is at https://anonymous.4open.science/r/SHNN. More details are in Appendix.
@@ 6.2 Comparison with Baselines
Table 3: Node classification accuracy (%) for HNNs and SHNN. The best result on each dataset is highlighted in bold font. The second and third highest accuracies are marked with an underline.
@@
Figure 2: The relative training time required for HNNs and SHNN to achieve optimal performance.
@@
Figure 3: Hyperlink prediction AUC (%) and relative training time for HNNs and SHNN.
Node classification. We summarize the classification accuracy and relative training time of the SHNN and HNNs in Table 3 and Figure 2, respectively. Table 3 demonstrates that SHNN not only leads to the best results across four datasets (Cora-CA, Congress, House, and Senate) but also results in the highest average mean accuracy overall. These findings confirm the effectiveness of SHNN in generating powerful node features for node classification. As discussed in Section 4, we attribute this effectiveness to SHNN’s ability to match the power of existing HNNs in utilizing neighbourhood information to generate node features, while requiring the optimization of fewer parameters. This makes the SHNN more efficient in utilizing training data and reduces the risk of overfitting, thereby outperforming more complex HNN counterparts. On the other hand, as shown in Figure 2, the SHNN consistently needs less training time to achieve superior performance compared to the HNNs. This underscores the high training efficiency of SHNN in downstream tasks.
Hyperlink prediction. The results for the hyperlink prediction task are summarized in Figure 3. These results exhibit a pattern similar to that observed in the hypergraph node classification task: the SHNN framework outperforms the HNNs while requiring less training time. An interesting observation is that the training efficiency improvement brought by SHNN is smaller in the hyperlink prediction task compared to the node classification task. The intuition behind this observation is that the task-specific block used for hyperlink prediction is more computationally expensive than the one used for hypergraph node classification. Specifically, the hyperlink predictor needs to generate a prediction for each hyperedge by aggregating the features of the nodes connected by the hyperedge, whereas the node classifier only needs to perform forward propagation for each individual node. We theoretically discuss that the efficiency improvement provided by SHNN is inversely correlated with the complexity of the task-specific block in Appendix.
@@ 6.3 Ablation study
We conduct a series of ablation studies on both the SPM-Block and the general SHNN model in node classification on the Cora-CA, Citeseer, House, and Senate datasets.
@@
Figure 4: The ablation study on the hyperparameters of SPM-Block in node classification.
Table 4: The ablation study on the components of SHNN in node classification. The best result on each dataset is highlighted in bold font.
SMP-Block Weighted $\mathbf{S}$ Cora-CA Citeseer Senate House Avg. Mean
@@
Ablation study on the hyperparameters of SPM-Block. We focus on two key hyperparameters of the SPM-Block: the number of layers and the $\alpha$. As illustrated in Figure 4, models with $\alpha>0$ exhibit more consistent accuracy as the number of layers increases compared to models with $\alpha=0$. We attribute this consistency to the robustness discussed in Proposition 4.2, where a positive $\alpha$ helps retain the distinct information of each node, thereby mitigating the oversmoothing issue. Additionally, we observe that a positive $\alpha$ significantly enhances performance in the House and Senate datasets, which are more heterophilic, compared to the Cora-CA and Citeseer datasets. This suggests that retaining the distinct information of each node is more beneficial for heterophilic datasets.
Ablation study on the components of SHNN. In this study, we examine the influence of the SMP-Block and the hyperedge-size-based clique expansion on the performance of the SHNN in node classification. Without the SMP-Block, the model functions as a pure MLP. We define $\mathbf{S}$ based on hyperedge-size-based clique expansion as weighted $\mathbf{S}$; in contrast, for the framework without weighted $\mathbf{S}$, $\mathbf{S}$ is generated using a $\mathbf{W}$ with all positive values set to one. As shown in Table 4, the inclusion of SMP-Block improves classification accuracy across all datasets, demonstrating that SMP-Block effectively incorporates hypergraph structural information into node features. Furthermore, adding weighted $\mathbf{S}$ further enhances performance, highlighting the informativeness of the hyperedge-size-based clique expansion. Thus, the components we designed significantly aid SHNN in leveraging useful information from the given hypergraph structure.
@@ 7 Conclusion & Limitation
In this paper, we propose a novel model called the simplified hypergraph neural network (SHNN). The key innovation of SHNN is an original training-free messgea passing block tailored for data on hypergraphs. We present both theoretical and empirical results demonstrating that SHNN can solve hypergraph-related downstream tasks with remarkable efficiency and effectiveness. To our knowledge, SHNN is the first model to integrate the hypergraph structural information in the data preprocessing stage, which significantly enhances training efficiency. The proposed SHNN can advance the field of hypergraph machine learning research by not only improving the efficiency of using observed hypergraph structures to solve downstream tasks, but also serving as a simple starting point for the development of future hypergraph machine learning models.
Limitation and future work. Learning tasks on hypergraphs can generally be defined at the node, hyperedge, and hypergraph levels. In this paper, we empirically test our SHNN on node-level and hyperedge-level tasks, with the main limitation being the lack of empirical results on hypergraph-level tasks. To our knowledge, designing task-specific blocks to generate hypergraph-level features based on existing node features itself remains a challenging topic in the literature. Therefore, we consider the application of our SHNN to hypergraph-level tasks as future work.